OpenStudy (anonymous):

The smaller of two similar rectangles has dimensions 4 and 6. Find the dimensions of the larger rectangle if the ratio of the perimeters is 2 to 3. 2 2/3 by 4 6 by 9 8 by 18 12 by 12

OpenStudy (anonymous):

The perimeter of the smaller rectangle is 2(6)+2(4) = 20 The ratio between the two rectangles is 2:3 2/3 = 20/x Solve the ratio and you have the perimeter of the larger rectangle.

OpenStudy (anonymous):

oh okay I get it xD thanks

OpenStudy (anonymous):

Perimeter = 30...dimensions will be 2:3 6/x = 2/3 4/x = 2/3 solve these ratios and you'll have the dimensions

OpenStudy (anonymous):

okay wait now I'm lost xD I got x=30